# Millimeter emission from supermassive black hole coronae

### v1

@@ 5 Conclusions
We presented a comprehensive study of mm-wave coronal emission in SMBH coronae. We first introduced a model for calculating the synchrotron emission from a SMBH corona based on Inoue & Doi (2018), coupled with an SED fitting scheme optimised for realistic datasets. That is, we minimised the number of free parameters in the model by introducing a physically motivated parameterization of the magnetic field intensity and an empirical parameterization for the opacity and temperature of the corona. We applied our model systematically to the five RQ AGNs for which coronal mm emission had been reported, together with two additional RQ AGNs with excellent multi-wavelength data from radio-cm to sub-mm frequencies. Our SED fitting results suggest that the coronae have non-thermal particle energy density fractions of 0.5–10% and radii of 60–250 $R_{\mathrm{g}}$, from which we infer typical magnetic fields of 10–60 G. These sets of parameters can also explain the putative correlation between $M_{\mathrm{BH}}$ and mm-wave luminosity presented in Ruffa et al. (2024). This opens the possibility of estimating $M_{\mathrm{BH}}$ via mm-wave observations in systems difficult to observe at other wavelengths, such as extremely obscured sources. Another suggested application of this approach is to measure faint coronal emission in high-$z$ lensed quasars using next-generation radio observatories. We conclude that multifrequency continuum observations spanning from radio to far-IR frequencies—with particular emphasis on the mm band—can be used as direct probes of the physical properties of SMBH coronae. The knowledge gained from these observations can have direct applications to other studies related to accretion disc coronae, such as in XRBs and tidal disruption events. Moreover, the detailed information on electron acceleration in the corona can help multi-messenger emission modelling, allowing to derive more robust predictions of neutrino sources, or verify neutrino detections.
@@ Acknowledgements.
SdP, CY, SA and SK gratefully acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 789410, PI: S. Aalto). SGB acknowledges support from the Spanish grant PID2022-138560NB-I00, funded by MCIN AEI/10.13039/501100011033/FEDER, EU. MPS acknowledges support under grants RYC2021-033094-I, CNS2023-145506 and PID2023-146667NB-I00 funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. JBT acknowledges support from the DFG, via the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters - from Source to Signal” (project no. 445052434). IGB is supported by the Programa Atracción de Talento Investigador “César Nombela” via grant 2023-T1/TEC-29030 funded by the Community of Madrid. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2011.0.01234.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. This research has made use of the CIRADA cutout service at cutouts.cirada.ca, operated by the Canadian Initiative for Radio Astronomy Data Analysis (CIRADA). CIRADA is funded by a grant from the Canada Foundation for Innovation 2017 Innovation Fund (Project 35999), as well as by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec, in collaboration with the National Research Council of Canada, the US National Radio Astronomy Observatory and Australia’s Commonwealth Scientific and Industrial Research Organisation.
@@ Appendix A Synchrotron emission from a thermal population of electrons
We explore in more detail whether it is feasible for thermal electrons alone to produce appreciable synchrotron emission, compatible with the mm emission detected in several sources. The number of relativistic electrons in a thermal distribution strongly depends on the temperature of the plasma. We, therefore, calculated SEDs for corona temperatures ranging between 200–400 keV, shown in Fig. 9. We note that values of $kT_{\mathrm{c}}\sim 500$ keV could be unsustainable due to runaway pair production. For $kT_{\mathrm{c}}<200$ keV, there are simply not enough relativistic electrons to emit detectable synchrotron, and even for $kT_{\mathrm{c}}>300$ keV this emission is faint. For reference, we also show the total emission, including a very small fraction of non-thermal electrons, $\delta=10^{-3}$. The non-thermal electrons completely dominate the emission for $kT_{\mathrm{c}}<300$ keV. The only case in which the emission from thermal electrons could be dominant is for extremely high corona temperatures $kT_{\mathrm{c}}\sim 400$ keV and low values of $\delta\lesssim 10^{-3}$. However, such a scenario ($kT_{\mathrm{c}}\sim 400$ keV, $\tau_{\mathrm{T}}\sim 1$, $r_{\mathrm{c}}\sim 120$, $B\sim 100$ G) is energetically very demanding, still produces faint mm emission, and would most likely overpredict the X-ray luminosity. We thus conclude that scenarios with $kT_{\mathrm{c}}<200$ keV and $\delta>10^{-3}$ are far more promising for explaining the mm emission in RQ AGNs.
@@ Appendix B Source sample
In the following, we provide a short description of the sources analysed in this work and their SED fitting. The fluxes reported in this work are a combination of archival and proprietary observations. Among the archives used, we highlight the LoTSS survey with LOFAR (Shimwell et al. 2017), the RACS survey with ASKAP (Hale et al. 2021), and the VLASS survey with the VLA (Lacy et al. 2020). We note that all flux density errors presented in the tables include only the statistical errors; for the inclusion of systematic errors, we refer to Sect. 2.2. For flux densities that are not taken from the literature, we also specify the observatory that took the data we used. We used CARTA [^9] https://cartavis.org/ to obtain the flux density fitting a Gaussian profile.
@@ B.1 GRS 1734--292
This galaxy is located at a distance of 84 Mpc ($z=0.021$) and is a known hard X-ray source (e.g. Tortosa et al. 2017). Moreover, Michiyama et al. (2024) found a $\gtrsim$50increase in the source 100-GHz flux in timescales of days; this mm flux variability is a strong indication of the origin of the mm flux in the corona (we refer to Shablovinskaya et al. 2024, for an in-depth discussion of this reasoning). Unfortunately, the source also presents a weak radio jet on angular scales of a few arcseconds, making it difficult to build a consistent SED, as the presence of diffuse emission leads to an artificially steep spectrum if the lower resolution data at lower frequencies are included. Due to these limitations on the dataset, and the fact that f–f emission cannot dominate at $\sim$100 GHz, we simply opt to remove the f–f component and to fix $\alpha_{\mathrm{sy}}=0.75$ as in Michiyama et al. (2024). In addition, the lack of data above 230 GHz prevents us from constraining the dust component, so we fix $\beta=1.78$ and $\nu_{\tau_{1}}=800$ GHz. We adopt a value of $M_{\mathrm{BH}}=10^{7.84}$ M (Mejía-Restrepo et al. 2022). Furthermore, we fix $\tau_{\mathrm{T}}=2.98$ and $kT_{\mathrm{c}}=12$ keV, as derived by Tortosa et al. (2017). We fit the flux densities listed in Table 2. We obtain $r_{\mathrm{c}}=112\pm 14$ ($R_{\mathrm{c}}\approx 0.5$ ld) and $\log{\delta}=-1.74\pm 0.13$, leading to $B\sim 21$ G and $\sigma\sim 0.011$; the corona peaks component at $\nu_{\mathrm{p}}=130\pm 14$ GHz with a peak flux density of $S_{\mathrm{p}}=0.91\pm 0.12$ mJy and has a luminosity of $L_{\mathrm{c}}\approx 4\times 10^{39}$ erg s-1. The predicted X-ray luminosity is $L_{\mathrm{14-150\,keV}}\approx 8\times 10^{43}$ erg s-1, corresponding to $\lambda_{\mathrm{Edd}}\sim 0.004$. The cornerplot of the fit is shown in Fig. 11.
@@
Table 2: Flux densities and sizes of the flux extraction regions for GRS 1734--292. The Type column indicates if the values were used for the fit, considered as upper limits (UL) or lower limits (LL). References: (a)Michiyama et al. (2024) and references therein; (b)this work; (c)Marti et al. (1998).
@@ B.2 IC 4329A
This is a very well-studied source both in radio and X-rays (e.g. Inoue & Doi 2018; Ingram et al. 2023; Tortosa et al. 2024; Shablovinskaya et al. 2024), showing convincing evidence of a hybrid corona. In fact, this source is one of the first two sources for which a bump in the SED around $\sim$100 GHz was detected (Inoue & Doi 2018). Moreover, Shablovinskaya et al. (2024) proved that the emission at 100 GHz is significantly variable on daily timescales. We adopt $M_{\mathrm{BH}}=6.8\times 10^{7}$  M as in Shablovinskaya et al. (2024), and fit our model to the same dataset presented in Table 1 of Inoue & Doi (2018). To present a consistent approach as with the other sources, we fix $p=2.7$ and $\tau_{\mathrm{T}}=0.25$ and $kT_{\mathrm{c}}\approx 166$ K. We note that, observationally, the values of $\tau_{\mathrm{T}}$ and $kT_{\mathrm{c}}$ are not well constrained and are also variable (e.g. Tortosa et al. 2024), but the specific values adopted should not play a significant role in the fitting (as shown in Sect. 3.1). The SED is poorly constrained between 10 and 100 GHz; we therefore fix $\alpha_{\mathrm{sy}}=-0.59$ as in Inoue & Doi (2018). We obtain $r_{\mathrm{c}}=246^{+90}_{-36}$ and $\log{\delta}=-2.15^{+0.17}_{-0.38}$ ($R_{\mathrm{c}}\approx 1$ ld), leading to $B\sim 10.6$ G and $\sigma\sim 0.076$. The corona component peaks at $\nu_{\mathrm{p}}=95\pm 13$ GHz, with a peak flux density of $S_{\mathrm{p}}=4\pm 1$ mJy, and has a luminosity of $L_{\mathrm{c}}\approx 8\times 10^{39}$ erg s-1. The cornerplot of the fit is shown in Fig. 12. The predicted X-ray luminosity is $L_{\mathrm{14-150\,keV}}\approx 2\times 10^{44}$ erg s-1, corresponding to $\lambda_{\mathrm{Edd}}\sim 0.02$. Alternatively, the SED fit to the multiple epochs of 100-GHz observations presented in Shablovinskaya et al. (2024) yielded a broader range of sizes of the corona of $\sim$170–300 $R_{\mathrm{g}}$.
Table 3: Flux densities and sizes of the flux extraction regions for IC 4329A. The Type column indicates if the values were used for the fit, which are considered upper limits (UL) or lower limits (LL). References: (a)Inoue & Doi (2018); (b)Ichikawa et al. (2019); (c)this work.
@@ B.3 MCG--06--30--15
This RQ-AGN is located at a distance of $D=34$ Mpc and hosts a SMBH with $\log{M_{\mathrm{BH}}}\approx 7.3$ (Raimundo et al. 2013, and references therein). From hard X-ray observations it is inferred $kT_{\mathrm{c}}\approx 63$ keV and $\tau_{\mathrm{T}}=0.27$. An analysis of the VLA fluxes from images at different resolutions indicates that there is significant diffuse emission on scales $>0.3"$ (Table 4).
@@
Table 4: Flux densities and sizes of the flux extraction regions for MCG--06--30--15. The Type column indicates if the values were used for the fit, considered as upper limits (UL) or lower limits (LL). References: (a)this work; (b)Ichikawa et al. (2019).
@@
90.52 $0.59\pm 0.02$ $0.090\times 0.090$ yes ALMA(a)
92.42 $0.59\pm 0.02$ $0.090\times 0.090$ yes ALMA(a)
102.52 $0.57\pm 0.02$ $0.090\times 0.090$ yes ALMA(a)
104.48 $0.60\pm 0.02$ $0.090\times 0.090$ yes ALMA(a)
226.31 $0.56\pm 0.02$ $0.123\times 0.123$ yes ALMA(a)
240.83 $0.53\pm 0.02$ $0.123\times 0.123$ yes ALMA(a)
228.81 $0.59\pm 0.02$ $0.123\times 0.123$ yes ALMA(a)
243.08 $0.54\pm 0.02$ $0.123\times 0.123$ yes ALMA(a)
@@ B.4 MCG+08--11--11
Table 5: Flux densities and sizes of the flux extraction regions for MCG+08--11--11. The Type column indicates if the values were used for the fit, considered as upper limits (UL) or lower limits (LL). References: (a)this work; (b)Smith et al. (2020); (c)Panessa et al. (2022); (d)Behar et al. (2018); (e)Petrucci et al. (2023); (f)Ichikawa et al. (2019).

### v2

@@ 5 Conclusions
We presented a comprehensive study of mm-wave coronal emission in SMBH coronae. We first introduced a model for calculating the synchrotron emission from an SMBH corona based on Inoue & Doi (2018), coupled with an SED fitting scheme optimised for realistic datasets. That is, we minimised the number of free parameters in the model by introducing a physically motivated parameterization of the magnetic field intensity and an empirical parameterization for the opacity and temperature of the corona. We applied our model systematically to the five RQ AGNs for which coronal mm emission had been reported, together with two additional RQ AGNs with excellent multi-wavelength data from radio-cm to sub-mm frequencies. Our SED fitting results suggest that the coronae have non-thermal particle energy density fractions of 0.5–10% and radii of 60–250 $R_{\mathrm{g}}$, from which we infer typical magnetic fields of 10–60 G. These sets of parameters can also explain the putative correlation between $M_{\mathrm{BH}}$ and mm-wave luminosity presented in Ruffa et al. (2024). This opens the possibility of estimating $M_{\mathrm{BH}}$ via mm-wave observations in systems difficult to observe at other wavelengths, such as extremely obscured sources. Another suggested application of this approach is to measure faint coronal emission in high-$z$ lensed quasars using next-generation radio observatories. We conclude that multifrequency continuum observations spanning from radio to far-IR frequencies—with particular emphasis on the mm band—can be used as direct probes of the physical properties of SMBH coronae. The knowledge gained from these observations can have direct applications to other studies related to accretion disc coronae, such as in XRBs and tidal disruption events. Moreover, the detailed information on electron acceleration in the corona can help multi-messenger emission modelling, allowing to derive more robust predictions of neutrino sources, or verify neutrino detections.
@@ Acknowledgements.
SdP, CY, SA and SK gratefully acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 789410, PI: S. Aalto). SGB acknowledges support from the Spanish grant PID2022-138560NB-I00, funded by MCIN AEI/10.13039/501100011033/FEDER, EU. MPS acknowledges support under grants RYC2021-033094-I, CNS2023-145506 and PID2023-146667NB-I00 funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. JBT acknowledges support from the DFG, via the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters - from Source to Signal” (project no. 445052434). IGB is supported by the Programa Atracción de Talento Investigador “César Nombela” via grant 2023-T1/TEC-29030 funded by the Community of Madrid. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2017.1.00236.S, ADS/JAO.ALMA#2019.1.00618.S, ADS/JAO.ALMA#2019.1.01230.S, ADS/JAO.ALMA#2015.1.01047.S, ADS/JAO.ALMA#2023.1.00121.S, ADS/JAO.ALMA#2021.1.00812.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. This research has made use of the CIRADA cutout service at cutouts.cirada.ca, operated by the Canadian Initiative for Radio Astronomy Data Analysis (CIRADA). CIRADA is funded by a grant from the Canada Foundation for Innovation 2017 Innovation Fund (Project 35999), as well as by the Provinces of Ontario, British Columbia, Alberta, Manitoba and Quebec, in collaboration with the National Research Council of Canada, the US National Radio Astronomy Observatory and Australia’s Commonwealth Scientific and Industrial Research Organisation.
@@ Appendix A Synchrotron emission from a thermal population of electrons
We explore in more detail whether it is feasible for thermal electrons alone to produce appreciable synchrotron emission, compatible with the mm emission detected in several sources. The number of relativistic electrons in a thermal distribution strongly depends on the temperature of the plasma. We, therefore, calculated SEDs for corona temperatures ranging from 200 to 400 keV, shown in Fig. 9. We note that values of $kT_{\mathrm{c}}\sim 500$ keV could be unsustainable due to runaway pair production. For $kT_{\mathrm{c}}<200$ keV, there are simply not enough relativistic electrons to emit detectable synchrotron, and even for $kT_{\mathrm{c}}\gtrsim 300$ keV this emission is quite faint. For reference, we also show the total emission, including a very small fraction of non-thermal electrons, $\delta=10^{-3}$. The non-thermal electrons completely dominate the emission for $kT_{\mathrm{c}}<300$ keV. The only case in which the emission from thermal electrons could be dominant is for extremely high corona temperatures $kT_{\mathrm{c}}\sim 400$ keV and low values of $\delta\lesssim 10^{-3}$. However, such a scenario ($kT_{\mathrm{c}}\sim 400$ keV, $\tau_{\mathrm{T}}\sim 1$, $r_{\mathrm{c}}\sim 120$, $B\sim 100$ G) is energetically very demanding, still produces faint mm emission, and would most likely overpredict the X-ray luminosity. We thus conclude that scenarios with $kT_{\mathrm{c}}<200$ keV and $\delta>10^{-3}$ are far more promising for explaining the mm emission in RQ AGNs.
@@ Appendix B Source sample
In the following, we provide a short description of the sources analysed in this work and their SED fitting. The fluxes reported in this work are a combination of archival and proprietary observations. As such, the datasets are non-simultaneous, with typical timespans of several years. Among the archives used, we highlight the LoTSS survey with LOFAR (Shimwell et al. 2017), the RACS survey with ASKAP (Hale et al. 2021), and the VLASS survey with the VLA (Lacy et al. 2020). We note that all flux density errors presented in the tables include only the statistical errors; for the inclusion of systematic errors, we refer to Sect. 2.2. For flux densities that are not taken from the literature, we also specify the observatory that took the data we used. We used CARTA [^10] https://cartavis.org/ to obtain the flux density fitting a Gaussian profile.
@@ B.1 GRS 1734--292
This galaxy is located at a distance of 84 Mpc ($z=0.021$) and is a known hard X-ray source (e.g. Tortosa et al. 2017). Moreover, Michiyama et al. (2024) found a $\gtrsim$50increase in the source 100-GHz flux in timescales of days; this mm flux variability is a strong indication of the origin of the mm flux in the corona (we refer to Shablovinskaya et al. 2024, for an in-depth discussion of this reasoning). Unfortunately, the source also presents a weak radio jet on angular scales of a few arcseconds, making it difficult to build a consistent SED, as the presence of diffuse emission leads to an artificially steep spectrum if the lower resolution data at lower frequencies are included. Due to these limitations on the dataset, and the fact that f–f emission cannot dominate at $\sim$100 GHz, we simply opt to remove the f–f component and to fix $\alpha_{\mathrm{sy}}=0.75$ as in Michiyama et al. (2024). In addition, the lack of data above 230 GHz prevents us from constraining the dust component, so we fix $\beta=1.78$ and $\nu_{\tau_{1}}=800$ GHz. We adopt a value of $M_{\mathrm{BH}}=10^{7.84}$ M (Mejía-Restrepo et al. 2022). Furthermore, we fix $\tau_{\mathrm{T}}=2.98$ and $kT_{\mathrm{c}}=12$ keV, as derived by Tortosa et al. (2017). We fit the flux densities listed in Table 2. We obtain $r_{\mathrm{c}}=112\pm 14$ ($R_{\mathrm{c}}\approx 0.5$ ld) and $\log{\delta}=-1.74\pm 0.13$, leading to $B\sim 20$ G and $\sigma\sim 0.011$; the corona peaks at $\nu_{\mathrm{p}}=130\pm 14$ GHz with a peak flux density of $S_{\mathrm{p}}=0.91\pm 0.11$ mJy and has a luminosity of $L_{\mathrm{c}}\approx 4\times 10^{39}$ erg s-1. The predicted X-ray luminosity is $L_{\mathrm{14-150\,keV}}\approx 8\times 10^{43}$ erg s-1, corresponding to $\lambda_{\mathrm{Edd}}\sim 0.004$. The cornerplot of the fit is shown in Fig. 11.
@@
Table 2: Flux densities and sizes of the flux extraction regions for GRS 1734--292. The Type column indicates if the values were used for the fit as observed (yes), considered as upper limits (UL) or lower limits (LL). References: (a)Michiyama et al. (2024) and references therein; (b)this work; (c)Marti et al. (1998).
@@ B.2 IC 4329A
This is a very well-studied source both in radio and X-rays (e.g. Inoue & Doi 2018; Ingram et al. 2023; Tortosa et al. 2024; Shablovinskaya et al. 2024), showing convincing evidence of a hybrid corona. In fact, this source is one of the first two sources for which a bump in the SED around $\sim$100 GHz was detected (Inoue & Doi 2018). Moreover, Shablovinskaya et al. (2024) proved that the emission at 100 GHz is significantly variable on daily timescales. We adopt $M_{\mathrm{BH}}=6.8\times 10^{7}$  M as in Shablovinskaya et al. (2024), and fit our model to the same dataset presented in Table 1 of Inoue & Doi (2018). To present a consistent approach as with the other sources, we fix $p=2.7$ and $\tau_{\mathrm{T}}=0.25$ and $kT_{\mathrm{c}}\approx 166$ K. We note that, observationally, the values of $\tau_{\mathrm{T}}$ and $kT_{\mathrm{c}}$ are not well constrained and are also variable (e.g. Tortosa et al. 2024), but the specific values adopted should not play a significant role in the fitting (as shown in Sect. 3.1). The SED is poorly constrained between 10 and 100 GHz; we therefore fix $\alpha_{\mathrm{sy}}=-0.59$ as in Inoue & Doi (2018). We obtain $r_{\mathrm{c}}=261^{+141}_{-45}$ and $\log{\delta}=-2.25^{+0.24}_{-0.48}$ ($R_{\mathrm{c}}\approx 1$ ld), leading to $B\sim 9.3$ G and $\sigma\sim 0.06$. The corona component peaks at $\nu_{\mathrm{p}}=90\pm 20$ GHz, with a peak flux density of $S_{\mathrm{p}}=4.2\pm 1.1$ mJy, and has a luminosity of $L_{\mathrm{c}}\approx 7\times 10^{39}$ erg s-1. The cornerplot of the fit is shown in Fig. 12. The predicted X-ray luminosity is $L_{\mathrm{14-150\,keV}}\approx 2\times 10^{44}$ erg s-1, corresponding to $\lambda_{\mathrm{Edd}}\sim 0.02$. Alternatively, the SED fit to the multiple epochs of 100-GHz observations presented in Shablovinskaya et al. (2024) yielded a broader range of sizes of the corona of $\sim$170–300 $R_{\mathrm{g}}$.
Table 3: Flux densities and sizes of the flux extraction regions for IC 4329A. The Type column indicates if the values were used for the fit as observed (yes), which are considered upper limits (UL) or lower limits (LL). References: (a)Inoue & Doi (2018); (b)Ichikawa et al. (2019); (c)this work; (d)(Imanishi et al. 2016).
@@ B.3 MCG--06--30--15
This RQ-AGN is located at a distance of $D=34$ Mpc and hosts an SMBH with $\log{M_{\mathrm{BH}}}\approx 7.3$ (Raimundo et al. 2013, and references therein). From hard X-ray observations it is inferred $kT_{\mathrm{c}}\approx 63$ keV and $\tau_{\mathrm{T}}=0.27$. An analysis of the VLA fluxes from images at different resolutions indicates that there is significant diffuse emission on scales $>0.3"$ (Table 4).
@@
Table 4: Flux densities and sizes of the flux extraction regions for MCG--06--30--15. The Type column indicates if the values were used for the fit as observed (yes), considered as upper limits (UL) or lower limits (LL). References: (a)this work; (b)Ichikawa et al. (2019).
@@
90.52 $0.59\pm 0.03$ $0.090\times 0.090$ yes ALMA(a)
92.42 $0.59\pm 0.03$ $0.090\times 0.090$ yes ALMA(a)
102.52 $0.57\pm 0.03$ $0.090\times 0.090$ yes ALMA(a)
104.48 $0.60\pm 0.03$ $0.090\times 0.090$ yes ALMA(a)
226.31 $0.56\pm 0.03$ $0.123\times 0.123$ yes ALMA(a)
240.83 $0.53\pm 0.03$ $0.123\times 0.123$ yes ALMA(a)
228.81 $0.59\pm 0.03$ $0.123\times 0.123$ yes ALMA(a)
243.08 $0.54\pm 0.03$ $0.123\times 0.123$ yes ALMA(a)
348.4 $1.0\pm 0.2$ $0.27\times 0.24$ UL ALMA(a)
@@ B.4 MCG+08--11--11
Table 5: Flux densities and sizes of the flux extraction regions for MCG+08--11--11. The Type column indicates if the values were used for the fit as observed (yes), considered as upper limits (UL) or lower limits (LL). References: (a)this work; (b)Smith et al. (2020); (c)Panessa et al. (2022); (d)Behar et al. (2018); (e)Petrucci et al. (2023); (f)Ichikawa et al. (2019).